# Dissecting Mahalanobis: How Feature Geometry and Normalization Shape OOD Detection

### v1

@@ 4 Geometry of Representations
In the previous section we showed that no single OOD method yield consistent performance and behavior across multiple models. In fact, different SSL models and pretraining regimes produce representations with distinct geometric properties, indicating that OOD performance depends on the intrinsic structure of the representation space. To understand these effects, we analyze the internal geometry of model representations, seeking to answer: What internal characteristics of a model’s feature space predict strong OOD detection, and how do pretraining and fine-tuning shape these characteristics?
4.1 Spectral analysis
To understand how the intrinsic geometry of representations affects OOD performance, we begin by examining the spectral properties of three key matrices: the feature covariance $C$, the within-class scatter $S_{w}$, and the between-class scatter $S_{b}$. These matrices capture complementary aspects of the feature space: $C$ reflects overall variance, $S_{w}$ measures intra-class dispersion, and $S_{b}$ quantifies inter-class separation (more details in Appendix A). Our first analysis focuses on the eigenvalue spectra of these matrices. The magnitude and decay of eigenvalues reveal how variance is distributed across dimensions, providing insight into the richness and anisotropy of the feature space. For instance, a steep decay in $S_{w}$ eigenvalues indicates that intra-class variability is concentrated along a few directions, resulting in tight clusters, whereas a slower decay suggests more diffuse intra-class variation. Similarly, large eigenvalues in $S_{b}$ correspond to well-separated class means, signaling strong discriminability.
Spectral ratios
Eigenvalue shifts
Beyond static spectra, we are interested in how stable the representation geometry is under distributional shifts. To capture this, we define a spectral shift metric, which measures the relative change in eigenvalues from the training set to validation or OOD data (see Appendix A). A small shift indicates that the representation preserves its structure across data splits, signaling robustness. Large positive shifts reveal that features are spreading along new directions, while large negative shifts indicate compression. Figure 4 shows that OOD samples induce larger spectral shifts in models trained on small datasets, reflecting lower generalization and brittle feature structures. Large-scale pretrained models show smaller shifts, indicating more stable, robust representations under distributional change. Fine-tuning generally maintains small shifts while increasing $S_{b}$, improving class separation without compromising intra-class compactness.
4.2 Geometric trade-offs
To systematically identify what makes a representation “good” for OOD detection, we correlate spectral and manifold-based metrics with detection performance. The geometry of the representations is assessed with two complementary families of measures: manifold-geometry metrics—including Intrinsic Dimensionality (ID) and Local Intrinsic Dimensionality (LID) (Ma et al., 2018)—and eigenvalue-based metrics (e.g., Entropy, Slope, Fisher Ratio) computed on the covariance matrix $C$ and the Fisher scatter matrices $S_{w}$ and $S_{b}$. A complete description of all metrics is provided in Appendix B. Correlation analysis, visualized in Figure 5, highlights several metrics that correlate strongly with OOD performance.
Figure 5: Spearman correlations between representation metrics and OOD performance across Mahalanobis variants. The three Mahalanobis-based detectors exploit different geometric cues, which explains their distinct correlation patterns. These correlations also hold for Pearson correlations (see Appendix F).
Key insights
Mahalanobis distance shows moderate correlation with individual metrics; its sensitivity to the global eigen-spectrum can make it less reliable for ambiguous OOD regions. Relative Mahalanobis emphasizes how well features fit their class cluster by normalizing out global variance. It correlates strongly with $S_{w}$ metrics, such as entropy and eigenvalue decay, reflecting the importance of compact, well-separated clusters. Marginal Mahalanobis ignores class structure and correlates primarily with global metrics ($C$ and $S_{b}$), indicating that its success depends on overall manifold shape rather than per-class separation.
4.3 Ideal geometry
OOD detection performance depends on both the magnitude of separation between ID and OOD features and the internal structure of the representation. Ideally, an effective representation exhibits a balance between local manifold complexity and intra-class compactness: low-dimensional manifolds require tighter clusters to separate OOD data, while high-dimensional manifolds allow looser clusters, as extra directions naturally push OOD samples away. This “ideal geometry” reflects a compensatory relationship between these two principles. However, as we observed in Sections 3.3 and 4.2, single metrics fail to satisfy both principles simultaneously. Models with larger raw feature separation can perform worse than models with smaller separation, indicating that additional geometric properties beyond simple separation are critical for reliable detection. While individual metrics provide insight, they capture only a single aspect of the geometry.
Figure 6: Correlation of the product of LID and $S_{w}$ slope with OOD detection performance. Minimizing the absolute value indicates an optimal balance between manifold richness and cluster compactness.
To address this limitation, we consider combinations of metrics across different geometric aspects. In particular, the product of LID and the slope of $S_{w}$ strongly predicts OOD performance, which is shown in Figure 6. For the standard MD, this combined metric achieves a Spearman correlation of $\rho_{s}=0.85$, and it remains highly predictive for RMD and MMD ($\rho_{s}=0.72$ and $0.74$, respectively).
In practice, when the feature manifold is locally simple (low LID), the model has fewer dimensions to separate ID and OOD data, requiring extremely tight and compact class clusters (a steep, negative slope). The detector thus relies on strict confinement of ID data to identify outliers. Conversely, when the manifold is locally rich and complex (high LID), the space itself helps isolate OOD samples, allowing less compact clusters (a shallower slope), as high dimensionality drives OOD separation. The product of LID and slope reaches its ideal when minimized in absolute magnitude, reflecting the optimal balance between local manifold richness and cluster compactness.
5 Normalization Effects
Our preceding analysis has established that the efficacy of Mahalanobis-based OOD detectors is intrinsically linked to the geometric structure of a model’s feature space. Yet, even when the geometry is favorable, standard Mahalanobis distance assumes Gaussian-distributed features with tied covariance, an assumption often violated in practice. Empirical studies, such as Mahalanobis++ (Mueller & Hein, 2025), show that deep neural network features frequently exhibit heavy-tailed distributions and large variations in feature norms. These deviations imply that ID features do not occupy a single, globally Euclidean space; instead, they lie on a collection of low-dimensional, non-Euclidean submanifolds. Consequently, the raw Mahalanobis metric can misestimate distances, undermining OOD detection. These observations motivate explicit control over feature geometry: by adjusting feature magnitudes we can either better satisfy the Mahalanobis assumptions or deliberately reshape the space to improve OOD separation.
Table 1: False-positive rate (FPR) across models (averaged over datasets) using different Mahalanobis variants: MD* uses the empirically optimal $\beta$, $\hat{\textbf{MD}}$ uses the regression-predicted $\hat{\beta}$, MD (standard) fixes $\beta=0$, and MD++ (Mahalanobis++) fixes $\beta=1$. The regression-guided detector generally outperforms fixed-$\beta$ settings.
Detector BEiTV2 In1k BEiTV2 In21k ViT ViT In21K ViT In21K In1k ViT-L In21K In1k DeiT3 DeiT3 In21k In1k DeiT3-L In22k In1k EVA02 EVA02 In21k EVA02 In21k In1k EVA02-L In22k In1k ViT CLIP In1k ViT CLIP In12k In1k Average MD* 0.365 0.244 0.445 0.493 0.327 0.229 0.425 0.348 0.320 0.378 0.468 0.369 0.351 0.364 0.246 0.358 $\hat{\text{MD}}$ 0.375 0.274 0.456 0.518 0.346 0.239 0.428 0.356 0.333 0.401 0.498 0.383 0.368 0.386 0.254 0.375 MD 0.402 0.436 0.457 0.532 0.357 0.253 0.433 0.376 0.366 0.534 0.565 0.408 0.370 0.402 0.335 0.415 MD++ 0.376 0.298 0.454 0.577 0.387 0.282 0.430 0.356 0.342 0.446 0.506 0.382 0.378 0.382 0.278 0.392 RMD* 0.363 0.312 0.442 0.378 0.365 0.259 0.386 0.331 0.335 0.426 0.358 0.373 0.308 0.366 0.294 0.353 $\hat{\text{RMD}}$ 0.367 0.324 0.443 0.392 0.375 0.268 0.390 0.338 0.347 0.442 0.368 0.383 0.314 0.370 0.297 0.361 RMD 0.391 0.331 0.449 0.398 0.376 0.269 0.408 0.366 0.376 0.440 0.367 0.403 0.326 0.403 0.325 0.375 RMD++ 0.373 0.325 0.446 0.398 0.376 0.269 0.399 0.351 0.359 0.440 0.377 0.391 0.318 0.386 0.309 0.368
5.1 Radially Scaled $\ell_{2}$ Normalization
To control the geometry of the representation space and mitigate sensitivity to raw feature norms, we introduce a tunable radial $\ell_{2}$ normalization that applies a radially symmetric map to each feature vector. For a feature $z\in\mathbb{R}^{d}\setminus\{0\}$, the transformation is defined as:
@@ 5.1 Radially Scaled $\ell_{2}$ Normalization
where the scalar parameter $\beta\in\mathbb{R}$ governs the amount of radial scaling. This mapping transforms the feature space by altering the radial distances from the origin while preserving the direction of each vector. The new radius of a feature is given by $\|\phi_{\beta}(z)\|=\|z\|^{1-\beta}$. This transformation induces a new Riemannian metric on the feature space (see Appendix C for details).
Figure 7: Distribution of empirically optimal $\beta$ for MD and RMD detectors across OOD datasets. High variability indicates model- and dataset-specific tuning is needed.
Intuitively, adjusting $\beta$ allows us to compress or expand the feature space radially, controlling how tightly features cluster near the origin or how much “empty space” surrounds them:
  • $\beta<0$: Radii are expanded, aggressively emphasizing differences between norms.
  • $\beta=0$: Radii are unchanged, corresponding to the standard Mahalanobis distance.
  • $0<\beta<1$: Radii are compressed, while preserving relative ordering.
  • $\beta=1$: All points are projected onto the unit sphere, equivalent to the MD++ normalization.
  • $\beta>1$: Features farther from the origin are pulled inward more strongly than those closer, creating strong radial contraction.
By tuning $\beta$, we obtain a flexible mechanism to shape the feature geometry, providing a simple yet powerful knob for improving OOD separation while maintaining compact ID clusters.
@@ 5.2 Predicting the Optimal $\beta$
Our experiments show that the optimal radial parameter $\beta$ is highly model- and dataset-dependent, reflecting the intrinsic geometry of the learned representations. Moderate positive values often align the feature distribution with the Gaussian, tied-covariance assumptions of the Mahalanobis detector, yet in some cases larger or even negative values yield stronger in/out-of-distribution separation. Consequently, a fixed choice of $\beta$ is rarely optimal. Figure 7 shows the empirically optimal $\beta$ values (searched over [2,3][-2,3] in $0.25$ steps) for MD and RMD detectors across different OOD datasets. The wide spread of optimal values underscores that a one-size-fits-all approach is ineffective.
Regression Framework.
To eliminate the need for tuning on the target OOD dataset, we train a regression model that predicts $\beta$ using in-distribution geometry metrics, while allowing it to learn from other OOD datasets. We adopt a Leave-One-Dataset-Out scheme: for each target OOD dataset, the regression model is trained on all other OOD datasets and their corresponding ID features, ensuring that the target OOD samples are never seen during training. This setup is conceptually similar to outlier exposure (Hendrycks et al., 2018), where access to auxiliary OOD data helps guide the detector, but crucially, here the model generalizes to completely unseen OOD distributions. Candidate predictors include spectral properties of the feature covariance, intrinsic dimensionality estimates, and other representation-geometry statistics; highly collinear features ($\rho>0.9$) are removed. The target variable is the $\beta$ value minimizing the false-positive rate (FPR) for each model–dataset pair. Let $\hat{\beta}$ denote the predicted value.
Figure 8: Predicted vs. optimal $\beta$ for MD under Leave-One-Dataset-Out validation. The diagonal indicates perfect prediction.
Results.
Table 1 reports FPR across models and datasets. The regression-predicted $\hat{\beta}$ consistently improves OOD detection compared to fixed baselines ($\beta=0$ for standard MD, $\beta=1$ for MD++), for both MD and RMD detectors. For MD, the regression achieves a MAE of $0.72$ and $R^{2}=0.25$ in predicting the optimal $\beta$; for RMD, the MAE is $0.89$ with $R^{2}=0.47$. While the regression does not perfectly recover the empirically optimal $\beta$, it captures sufficient geometric information from in-distribution features to meaningfully improve detection. Figure 8 visualizes predicted vs. optimal $\beta$ for MD. Points close to the diagonal indicate that the regression captures key trends even for previously unseen OOD datasets. This demonstrates a practical path for tuning radial $\ell_{2}$ scaling without access to target OOD data, leveraging ID feature structure and prior knowledge from other auxiliary OOD datasets. We present the complete results and comparisons with baseline OOD detectors in Appendix D.2.

### v3

@@ 5 Geometry of Representations
In the previous section, we showed that no single OOD method yields consistent performance and behavior across multiple models. In fact, different SSL models and pretraining regimes produce representations with distinct geometric properties, indicating that OOD performance depends on the intrinsic structure of the representation space. To understand these effects, we analyze the internal geometry of model representations, seeking to answer: What internal characteristics of a model’s feature space predict strong OOD detection?
5.1 Geometry-detector alignment
To identify which representation properties matter for OOD detection, we correlate detection performance with two complementary families of metrics: (i) manifold metrics, such as intrinsic dimensionality  (Ma et al., 2018), computed from the ID features $F$, and (ii) spectral metrics computed from the eigenspectra of the global covariance $C$ and Fisher scatter matrices $S_{w},S_{b}$ (see Appendix A for definitions). Figure 4 summarizes Spearman correlations across detectors and models.
As we can see, the RMD correlates most strongly with within-class geometry ($S_{w}$), reflecting the importance of compact, well-structured class clusters. MMD correlates primarily with global geometry ($C$ and $S_{b}$), indicating dependence on the overall manifold shape. Standard MD sits between these extremes, combining sensitivity to both cluster structure and the global eigenspectrum. In Appendix C, we provide a spectral analysis that offers insight into how pretraining and fine-tuning shape these characteristics.
5.2 Ideal geometry: a compensatory trade-off
OOD detection depends not only on mean ID-OOD separation but also on how ID variability is organized. Two geometric factors repeatedly appear across models: local degrees of freedom (how many directions are explored in a neighborhood) and within-class concentration (how tightly class clusters concentrate around their means). Intuitively, if the local manifold is simple (low LID), reliable detection requires very compact clusters; if the local manifold is richer (high LID), OOD samples can deviate along many directions, allowing for less compact clusters. This yields a compensatory trade-off between local dimension and within-class concentration. Empirically, this trade-off is captured by the product of LID ($m$) and the within-class spectral slope ($|s|$) magnitude. Figure 5 shows that $m\cdot|s|$ (LID $\times$ slope of $S_{w}$ spectrum) strongly predicts Mahalanobis OOD performance across models and variants. We estimate the dataset-average LID $m$ using the kkNN LID estimator with $k=50$ and keep $k$ fixed across all experiments for consistency. The product $m|s|$ is expected to be stable to small variations in $k$ within the local regime (see Table 4 in Appendix E.4).
Taken together, these results suggest that neither local dimensionality nor within-class concentration alone is sufficient to explain Mahalanobis OOD behavior. Instead, strong performance is associated with representations that jointly balance these two properties, for which $m|s|$ provides a compact summary. This suggests that if we can move a representation along the $m\times|s|$ trade-off curve post-hoc, we may tune Mahalanobis behavior without changing the backbone. In the following section, we show how this summary can be traced and optimized by a simple post-hoc deformation of the feature space and later provide a mechanistic explanation for why $m|s|$ is predictive across Mahalanobis variants.
Figure 5: A simple predictor of Mahalanobis OOD performance. The product $m|s|$ (LID $\times$ within-class spectral-slope magnitude) correlates with Mahalanobis-based OOD detection across variants; lower values indicate better performance.
6 Radial scaling as a geometric control knob
The preceding section identifies $m|s|$ as a detector-invariant summary of ID geometry that predicts Mahalanobis OOD performance. We now introduce a simple post-hoc mechanism to modify this geometry without changing the backbone: a one-parameter family of direction-preserving radial deformations. This family generalizes standard $\ell_{2}$ normalization (unit-sphere projection) and provides a continuous knob for navigating a trajectory $\beta\mapsto(m(\beta),s(\beta))$ using ID data only.
6.1 Why radial normalization?
Mahalanobis-style detectors are quadratic forms that depend on a shared covariance estimate. In practice, feature norms can vary substantially across samples and models, and this radial variability can dominate the covariance fit and inflate score overlap between ID and OOD. Prior work has shown that $\ell_{2}$ normalizing features (projecting to the unit sphere) can stabilize Mahalanobis OOD detection by reducing norm-driven variation and improving the fit of quadratic scores (Mueller & Hein, 2025). We choose a direction-preserving radial family because it is the minimal intervention that changes norm-driven geometry while preserving angular class structure.
Table 1: OOD detection (FPR@95, $\downarrow$) averaged over five OpenOOD datasets. MD uses features as-is ($\beta=0$), MD++ applies $\ell_{2}$ normalization ($\beta=1$). RS-MD selects a per-model radial exponent $\hat{\beta}$ from ID data by selecting optimal geometric proxy $P(\beta)=m(\beta)\lvert s(\beta)\rvert$ over the search grid. RMD denotes the relative Mahalanobis variant, with analogous RMD++ ($\beta=1$) and RS-RMD (proxy-selected $\hat{\beta}$) settings. For each model (column), RS-MD is highlighted in light green when it improves over both MD and MD++; likewise, RS-RMD when it outperforms both RMD and RMD++. Best (lowest) results in each column are bolded.
[…]
6.2 Radially scaled Mahalanobis distance
Given a feature vector $z\in\mathbb{R}^{d}\setminus\{0\}$, we define the radial map
@@ 6.2 Radially scaled Mahalanobis distance
where $\beta\in\mathbb{R}$ controls radial contraction/expansion while preserving direction. The induced radius is $\|\phi_{\beta}(z)\|=\|z\|^{1-\beta}$. Thus, for $\beta>1$, the map contracts norms greater than 1 and expands norms smaller than 1 (pushing toward the unit sphere); $\beta<1$ has the opposite tendency (pushing away from the unit sphere); and $\beta<0$ expands the norms (see Appendix D for more details). This family contains key special cases: $\beta=0$ recovers the original geometry, and $\beta=1$ projects features onto the unit sphere. In practice, most feature norms are >1>1.
Definition.
We denote by RS-MD the Mahalanobis-distance detector applied to features after the transformation $\phi_{\beta}$ in Eq. 5. Thus, $\beta=0$ recovers standard MD, and $\beta=1$ corresponds to the $\ell_{2}$-normalized variant (MD++). We define RS-RMD analogously for the relative Mahalanobis variant evaluated on $\phi_{\beta}(z)$.
@@ 6.3 Predicting the Optimal $\beta$
What Changes Under $\phi_{\beta}$.
Applying $\phi_{\beta}$ modifies both (i) local neighborhood structure (and thus LID $m(\beta)$) and (ii) the spectrum of within-class scatter $S_{w}(\beta)$ (and thus the slope $s(\beta)$). Thus, changing $\beta$ systematically alters (i) neighborhood geometry (captured by $m(\beta)$) and (ii) within-class spectral decay (captured by $s(\beta)$), which in turn shifts the quadratic scores through the refit scatter estimate.
@@
Our experiments show that the optimal radial parameter $\beta$ is highly model- and dataset-dependent, reflecting the intrinsic geometry of the learned representations. Moderate positive values often align the feature distribution with the Gaussian, tied-covariance assumptions of the Mahalanobis detector; yet in some cases, larger or even negative values yield stronger in/out-of-distribution separation. Consequently, a fixed choice of $\beta$ is rarely optimal. Figure 6 shows the empirically optimal $\beta$ values (searched over [2,2][-2,2] in $0.25$ steps) for MD and RMD detectors across different OOD datasets. The wide spread of optimal values underscores that a one-size-fits-all approach is ineffective.
6.4 ID-only selection of $\beta$ via Geometric Proxy
Setup.
We study ID-only selection of $\beta$ on ImageNet-trained transformer backbones spanning several families (BEiT-v2, CLIP, EVA, and ViT variants with different pre-training and fine-tuning regimes). For each backbone and each $\beta$ on a discrete grid $\mathcal{B}\subset[-2,2]$, we estimate two quantities on a held-out portion of the ImageNet-1K training data: (i) the dataset-average local intrinsic dimensionality (LID) $m(\beta)$ using a k=50k{=}50 nearest-neighbor estimator (as in Section 5.2), and (ii) the within-class spectral slope $s(\beta)$ obtained from the eigenspectrum of the within-class scatter matrix $S_{w}(\beta)$ computed in the feature space induced by $\phi_{\beta}(z)$. OOD detection is then evaluated using the same datasets, protocol, and metrics as in Section 4.1; unless stated otherwise, the selected $\hat{\beta}$ is applied unchanged for that model on ImageNet (ID) and on all OOD benchmarks.
Figure 6: Empirically optimal $\beta$ varies across OOD settings. Distribution of the empirically optimal $\beta$ for MD and RMD detectors across OOD datasets. The wide spread highlights substantial model- and dataset-specific variation, indicating that $\beta$ typically requires tuning per setting.
Proxy definition
For each $\beta$, we compute (i) the dataset-average LID $m(\beta)$ and (ii) the within-class spectral slope $s(\beta)$ from the eigenspectrum of $S_{w}(\beta)$ formed on $\phi_{\beta}(z)$. We combine them into a single ID-only geometry summary,
$P(\beta)\triangleq m(\beta)\,|s(\beta)|,$ (6)
and evaluate $P(\beta)$ over the grid $\mathcal{B}\subset[-2,2]$,
Selecting $\hat{\beta}$ from the proxy curve
Across models, $P(\beta)=m(\beta)|s(\beta)|$ typically has an interior turning point (often inverted-U; occasionally U-shaped, e.g., ViTs). Since boundary optima can be artifacts of the finite search range, we select the most pronounced interior turning point: the interior grid value farthest from the endpoint baseline $P_{\mathrm{end}}=(P(\beta_{\min})+P(\beta_{\max}))/2$ (Appendix D.4). This recovers the interior maximum for inverted-U curves and the interior minimum for U-shaped curves; if no clear interior turning point appears on the grid, we use the Appendix fallback. In Section 3 connect these regimes to the instability functionals $\mathcal{I}(\beta)$ and $\widehat{\mathcal{I}}(\beta)$. Note that Section 5.2 reports an across-model correlation at $\beta=0$, which need not determine the within-model optimum along $\beta$.
Figure 7: Proxy feature–FPR trade-off across $\beta$ on NINCO (standard MD). Each panel corresponds to a model configuration. The blue curve reports the $\mathrm{FPR}@95$ as a function of $\beta$, while the red curve (right axis) shows the in-distribution proxy feature $P(\beta)\triangleq m(\beta)|s(\beta)|$. The star indicates the $\beta$ selected by the proxy, and the black x denotes the oracle $\beta$ that minimizes $\mathrm{FPR}@95$.
OOD performance summary
Table 1 reports $\mathrm{FPR}@95$ across backbones and OOD datasets. The proxy-selected $\hat{\beta}$ consistently improves OOD detection relative to fixed choices (e.g., $\beta=0$ for standard MD and $\beta=1$ for MD++), for both MD and RMD. Although the proxy does not always match the empirically optimal $\beta$, it captures enough ID geometric structure to yield meaningful gains in detection performance. Figure 8 further quantifies selection quality for standard MD by plotting the absolute error to the oracle choice $\beta$, i.e., $|\mathrm{FPR}(\hat{\beta})-\mathrm{FPR}(\beta)|$, across OOD benchmarks and the same set of models as in Table 1. The proxy achieves lower error than the $\beta=0$ and $\beta=1$ baselines across datasets, with strong improvements on NINCO, whose samples were verified to be free of ID contamination. The proxy also reduces worst-case behavior, as reflected by a lower upper tail (fewer/lower outliers) in the error distribution. Full results and comparisons against baseline OOD detectors are provided in Appendix F.2.
Figure 8: Absolute FPR error relative to the oracle $\beta^{*}$ (standard MD). Boxplots show $|\mathrm{FPR}(\hat{\beta})-\mathrm{FPR}(\beta^{*})|$ across OOD benchmarks. The proxy-selected $\hat{\beta}$ consistently achieves lower error than the fixed baselines $\beta{=}0$ and $\beta{=}1$ on every dataset.
7 A Unified Stability Lens for Mahalanobis Variants
The previous sections establish two empirical facts: (i) Mahalanobis-style OOD detection exhibits large cross-model variability, and (ii) the composite geometry summary $m(\beta)\,|s(\beta)|$ tracks performance across models and along $\beta$-trajectories. We now provide a mechanistic lens that explains why a low-dimensional ID-only geometry summary can be predictive while also clarifying its limitations.
7.1 Exact size–stretch factorization of quadratic scores
Let $\delta_{\beta}(z)\in\mathbb{R}^{d}$ denote the detector-specific centered deviation in $\phi_{\beta}$-space (e.g., class-conditional for MD, global for MMD), and let $\Sigma(\beta)\in\mathbb{R}^{d\times d}$ be the corresponding tied ID scatter/covariance used by the detector. We assume $\Sigma(\beta)\succ 0$ for all $\beta$ under consideration, so that $\Sigma(\beta)^{-1}$ exists. [^1] In practice, we enforce invertibility via standard regularization. For any Mahalanobis-type quadratic score
$S_{\beta}(z)\;=\;\delta_{\beta}(z)^{\top}\Sigma(\beta)^{-1}\delta_{\beta}(z),$ (7)
define the whitened stretch factor
$W_{\beta}(z)\;\triangleq\;\frac{\delta_{\beta}(z)^{\top}\Sigma(\beta)^{-1}\delta_{\beta}(z)}{\|\delta_{\beta}(z)\|^{2}},$ (8)
which is well-defined whenever $\|\delta_{\beta}(z)\|>0$. [^2] We assume $\|\delta_{\beta}(z)\|>0$ almost surely under $\mathcal{D}_{\mathrm{ID}}$; empirically, exact zeros occur with negligible frequency and are dropped. Then the score admits the exact factorization
$S_{\beta}(z)=\|\delta_{\beta}(z)\|^{2}\,W_{\beta}(z)$ (9)
and hence $\log S_{\beta}(z)=\log\|\delta_{\beta}(z)\|^{2}+\log W_{\beta}(z)$. This separates score variability into a size channel ($\|\delta\|^{2}$) and a stretch channel ($W$) that depends on alignment with the whitening geometry $\Sigma(\beta)^{-1}$.
Relative Mahalanobis Distance.
Eq. (7) covers detectors whose score is a single quadratic form (MD, MMD). RMD is a difference of two quadratic forms that generally use distinct scatter estimates, $S_{\mathrm{RMD},\beta}(z)=S^{(1)}_{\beta}(z)-S^{(2)}_{\beta}(z)$, and therefore does not admit a single shared $\Sigma(\beta)$ in general. Our stability lens applies to the RMD term-wise: each component $S^{(j)}_{\beta}(z)=\delta^{(j)}_{\beta}(z)^{\top}\Sigma^{(j)}(\beta)^{-1}\delta^{(j)}_{\beta}(z)$ admits the same factorization and channel decomposition, and RMD behavior additionally depends on interactions between the two terms (e.g., cancelation via cross-covariances). We provide the precise term-wise formulation in Appendix E.5.
7.2 ID instability decomposes into size, stretch, and compensation
To quantify how variable the ID scores are at a fixed operating point, we study the instability functional
$\mathcal{I}(\beta)\;\triangleq\;\operatorname{Var}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big[\log S_{\beta}(z)\big].$ (10)
Using Eq. (9) and the variance-covariance decomposition, $\mathcal{I}(\beta)$ decomposes exactly as
$\mathcal{I}(\beta)=A_{\delta}(\beta)+A_{W}(\beta)+2A_{\times}(\beta),$ (11)
where
$A_{\delta}(\beta)$ $\triangleq\operatorname{Var}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big[\log\|\delta_{\beta}(z)\|^{2}\big],$ (12)
$A_{W}(\beta)$ $\triangleq\operatorname{Var}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big[\log W_{\beta}(z)\big],$
$A_{\times}(\beta)$ $\triangleq\operatorname{Cov}_{z\sim\mathcal{D}_{\mathrm{ID}}}\!\big(\log\|\delta_{\beta}(z)\|^{2},\ \log W_{\beta}(z)\big).$
Empirically, both $A_{\delta}$ and $A_{W}$ vary substantially across models and across $\beta$, while $A_{\times}$ is typically negative, indicating systematic compensation between size and stretch (Appendix E.1). Consequently, no single scalar geometry statistic should be expected to mediate performance universally.
7.3 Stretch variability is driven by spectrum $\times$ allocation geometry
Let $\Sigma(\beta)=U(\beta)\Lambda(\beta)U(\beta)^{\top}$ and define the allocation of the deviation onto the eigenbasis by
$p_{i}(z;\beta)\triangleq\frac{(u_{i}(\beta)^{\top}\delta_{\beta}(z))^{2}}{\|\delta_{\beta}(z)\|^{2}},\qquad\sum_{i=1}^{d}p_{i}(z;\beta)=1.$ (13)
Then the stretch factor admits the exact identity
$W_{\beta}(z)=\sum_{i=1}^{d}\frac{p_{i}(z;\beta)}{\lambda_{i}(\beta)}.$ (14)
Equation (14) makes explicit that $A_{W}$ is controlled by the interaction of spectral heterogeneity ($\lambda_{i}(\beta)$) and pointwise allocation geometry ($p(z;\beta)$). We validate this mechanism with targeted interventions: rotations (changing allocations), allocation smoothing (collapsing allocation variability), and spectrum shaping (changing eigenvalues), all of which causally affect $A_{W}$ as predicted (Appendix E.3).
7.4 A compact ID-only proxy law along $\beta$
Motivated by Eq. (11)–(14), we test whether $\mathcal{I}(\beta)$ can be summarized by a small set of ID geometry statistics. Within each model, we find that the instability curve is well captured by
$\widehat{\mathcal{I}}(\beta)=a\,\log m_{k}(\beta)+b\,|s(\beta)|+c\,\mathrm{curv}(\beta),$ (15)
where $m_{k}(\beta)$ is the kkNN LID estimate and $(|s(\beta)|,\mathrm{curv}(\beta))$ summarizes the within-class spectrum (Appendix E.1). Figure 9 shows representative overlays; aggregate correlations and fitting details are reported in Appendix E.4. Our practical $\beta$-selection in Sec. 6.4 remains coefficient-free via $P(\beta)=m(\beta)|s(\beta)|$. Notably, the $\beta$-instability trend is model-specific: in Fig. 9, $\mathcal{I}(\beta)$ increases with $\beta$ for ViT, whereas it decreases for BEiTV2 and CLIP. These opposing trajectories explain why the preferred $\beta$ can lie in different regions across models and motivate treating distinct curve shapes as separate regimes tied to $\mathcal{I}(\beta)$ and $\widehat{\mathcal{I}}(\beta)$.
Figure 9: UST proxy tracks instability along $\beta$. For representative models, we plot $\mathcal{I}(\beta)=\operatorname{Var}(\log S_{\beta})$ (solid) and the ID-only proxy $\widehat{\mathcal{I}}(\beta)$ (dashed), both standardized within each model across $\beta$. The proxy tracks the shape of the instability trajectory and predicts its minimizer $\hat{\beta}$ using ID-only quantities. (Summary statistics across all models appear in Appendix E.4.)
Connection to $m(\beta)\,|s(\beta)|$.
Eq. (15) clarifies why $m(\beta)$ and $|s(\beta)|$ act as complementary coordinates: $|s|$ summarizes spectral heterogeneity that amplifies point-dependent exposure in Eq. (14), while $m_{k}$ captures an additional local geometry signal that improves the prediction of $A_{W}(\beta)$ beyond spectral statistics alone (see the coupling analysis in Appendix E.4). This provides a principled rationale for the empirical effectiveness of the simpler product $m(\beta)\,|s(\beta)|$ as a compact summary, while making explicit that it is necessarily lossy.